interest rate in India is equal 10%, and in USA it is 6%. What should be the exchange rate to equalize the interest rate. $1=70 INR
Answer #1

As per interest rate parity theory:

=1+rh/1+rf = F1/E0

=1.10/1.06 =F1/70

F1 = 72.6415